OpenStudy (anonymous):

a electric iron has rating of 750w;200v calculate 1 current 2 resistence of hating element

OpenStudy (radar):

Power in watts = EI = (E^2)/R = I^2(R) Use this and Ohm's Law. I will start you out. 750 = 200V* I I=750/200 I=3.75 amperes. For the resistance you can use Ohms Law, R=E/I = 200/3.75 A etc.